# Conformal geodesics are not variational in higher dimensions

### v1

@@ 1. Formulation of the Problem and the Result
Recently it was proven independently in [15] and [11] that in 3D the equation for un-parametrized conformal geodesics is variational; the corresponding Lagrangian is the torsion form. There exists a canonical family of parametrizations of conformal geodesics, so-called projective parameter [3]. Whatever parameter is used, the equation of parametrized conformal geodesics in 3D is however not variational [11].
The higher dimensional case has remained open. One could hope that a higher dimensional version of torsion functional (or any other conformally invariant Lagrangian) could solve the problem. That was verified in negative in [12] in the flat 4D case, using an invariant version of the variational bicomplex. The purpose of this note is to demonstrate the negative answer to this problem in full generality.
@@
Let $(M^{n},[g])$ be conformal manifold. If its dimension $n>3$, then the equation for conformal geodesics is not variational in the classical sense neither in parametrized nor in unparametrized form.
@@ 2. Warm-up: variational calculus and circles in three and two dimensions
Let us consider curves in the Euclidean space ${\mathbb{R}}^{3}(x,y,z)$. Unparametrized (generic) curves are given by equation of the type $y=y(x)$, $z=z(x)$, where $x$ can be seen as (technical) parameter, but which can be reparametrized. The curves have natural lifts to the jet-space $J^{3}({\mathbb{R}},{\mathbb{R}}^{2})={\mathbb{R}}^{9}(x,y_{0},z_{0},y_{1},z_{1},y_{2},z_{2},y_{3},z_{3})$ and the equation for conformal circles determines a submanifold in this space
@@
For a Lagrangian $L=L(x,y_{0},z_{0},y_{1},z_{1},y_{2},z_{2},\dots)$ the Euler $y$-operator is
$\frac{\delta L}{\delta y}=\frac{\partial L}{\partial y_{0}}-\frac{d}{dx}\Bigl{(}\frac{\partial L}{\partial y_{1}}\Bigr{)}+\frac{d^{2}}{dx^{2}}\Bigl{(}\frac{\partial L}{\partial y_{2}}\Bigr{)}-\dots$ (4)
and similar for $\frac{\delta L}{\delta z}$. The Euler-Lagrange equations EL are
@@
The inverse variational problem is to detect if the given equation, in our case (3), is of form (5) for some $L$. Let us stress that we do not require that the first equation is identical with the first Euler operator, etc, but the equations have to coincide as submanifolds in jets (inverse problem for differential operators is simpler and is solved via the so-called Helmholtz criterion [2]).
@@
The Lagrangian for variational equation of order $k$ can be chosen to be of order $<k$: this is achieved by modification of $L$ via divergence terms (integration by parts) and is known as Vainberg–Tonti method, see e.g. [13]. If $L$ is of order 2 then the coefficients of $y_{4},z_{4}$ in EL are $\frac{\partial^{2}L}{\partial y_{2}^{2}}$, $\frac{\partial^{2}L}{\partial y_{2}\partial z_{2}}$, $\frac{\partial^{2}L}{\partial z_{2}^{2}}$. Since (3) is of the third order, these expressions have to vanish and we have
$L=u(y_{0},z_{0},y_{1},z_{1})+v(y_{0},z_{0},y_{1},z_{1})y_{2}+w(y_{0},z_{0},y_{1},z_{1})z_{2}.$
Using the fact, that the Lagrangian is defined up to divergence, i.e. that the transformation $L\mapsto L+\frac{d}{dx}f$ does not change the Euler-Langrange equations, we can eliminate the coefficient $w\mapsto 0$. Then we set our equations (3) into EL and split it by 2-jets $y_{2},z_{2}$, receiving the following overdetermined PDE system:
$\displaystyle z_{1}u_{z_{0},y_{1}}=y_{1}^{2}v_{y_{0},y_{0}}+2y_{1}z_{1}v_{y_{0},y_{1}}+z_{1}^{2}v_{y_{1},y_{1}}-y_{1}u_{y_{0},z_{0}}+u_{y_{0}},\ z_{1}u_{y_{1},z_{1}}=-y_{1}u_{y_{0},z_{1}}+u_{y_{1}},$
$\displaystyle u_{z_{0},z_{1}}=v_{y_{1}},\ u_{z_{1},v}=0,\ y_{1}v_{y_{0},z_{1}}=-z_{1}v_{y_{1},z_{1}},\ z_{1}v_{z_{0},y_{1}}=-y_{1}v_{y_{0},z_{0}}+u_{z_{0},z_{0}}-2v_{y_{0}},$
$\displaystyle(1+y_{1}^{2}+z_{1}^{2})v_{z_{0},z_{1}}=-3y_{1}v_{z_{1}},\ (1+y_{1}^{2}+z_{1}^{2})v_{z_{1},z_{1}}=-3z_{1}v_{z_{1}}.$
@@
This Lagrangian slightly simplifies the one found in formula (11) of [11]. Prior to this a third order Lagrangian for conformal circles in Euclidean space was established in [4]. Note that only this latter Lagrangian (torsion: explored and generalized in [11]) is reparametrization-invariant.
@@
Considering the equation for parametrized circles
$x_{3}=-(x_{2}^{2}+y_{2}^{2})x_{1},\quad y_{3}=-(x_{2}^{2}+y_{2}^{2})y_{1}$
@@
$x_{3}=3\frac{x_{1}x_{2}+y_{1}y_{2}}{x_{1}^{2}+y_{1}^{2}}x_{2}-\frac{3}{2}\frac{x_{2}^{2}+y_{2}^{2}}{x_{1}^{2}+y_{1}^{2}}x_{1},\quad y_{3}=3\frac{x_{1}x_{2}+y_{1}y_{2}}{x_{1}^{2}+y_{1}^{2}}y_{2}-\frac{3}{2}\frac{x_{2}^{2}+y_{2}^{2}}{x_{1}^{2}+y_{1}^{2}}y_{1}$
we start with the Lagrangian $L=L(t,x_{0},y_{0},x_{1},y_{1},x_{2},y_{2})$ and then the condition that EL is of order $<4$ and simplification modulo divergences implies that we can take $L=u(t,x_{0},y_{0},x_{1},y_{1})+v(t,x_{0},y_{0},x_{1},y_{1})y_{2}$. However one of the EL equations is $\frac{\partial v}{\partial x_{1}}=0$, so $v=v(t,x_{0},y_{0},y_{1})$ and hence a change by a divergences can bring the Lagrangian to the form $L=u(t,x_{0},y_{0},x_{1},y_{1})$, for which the Euler-Lagrange equation is of order at most 2. This proves that in dimension 2 the analog equation for (conformal) circles (both in parametric and unparametric form) is not variational.
@@ 3. The inverse problem for conformal circles in four and five dimensions
Using the Vainberg–Tonti method, we can choose $L=L\bigl{(}x_{0}^{j},x_{1}^{j},x_{2}^{j}:1\leq j\leq 4\bigr{)}$. Denote the Euler-Lagrange operator
@@
$f^{1}_{3}f^{2}_{4}-f^{1}_{4}f^{2}_{3}+f^{1}_{4}f^{3}_{2}-f^{1}_{2}f^{3}_{4}+f^{1}_{2}f^{4}_{3}-f^{1}_{3}f^{4}_{2}+f^{2}_{1}f^{3}_{4}-f^{2}_{4}f^{3}_{1}+f^{2}_{3}f^{4}_{1}-f^{2}_{1}f^{4}_{3}+f^{3}_{1}f^{4}_{2}-f^{3}_{2}f^{4}_{1}$ (10)
which thus should be nonzero. Here we raise the indices $f^{i}=f_{i}$ using the Euclidean metric and denote $f^{i}_{j}$ the partial derivative of $f^{i}$ by $x^{j}_{1}$. The above expression is equivalent to nonvanishing of the following 4-form restricted to the graph $f^{j}=f^{j}(x^{j}_{1}:1\leq j\leq 4)$ where we frozen 0-jet variables $x^{j}_{0}$ and hid dependence on them; below we use $df^{j}=\sum_{i=1}^{4}f_{i}^{j}dx^{i}_{1}$, the indices $ijkl$ in the summation run over permutations of 12341234 12341234:
$\sum df^{i}\wedge df^{j}\wedge dx_{1}^{k}\wedge dx_{1}^{l}\neq 0.$ (11)
@@
$\displaystyle\frac{\partial^{2}f_{0}}{\partial u_{1}\partial u_{1}}=0,\quad 3u_{1}\frac{\partial f_{1}}{\partial u_{1}}+(x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2})\frac{\partial^{2}f_{1}}{\partial u_{1}\partial u_{1}}=0,\quad 3x_{1}\frac{\partial f_{1}}{\partial u_{1}}+(x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2})\frac{\partial^{2}f_{1}}{\partial x_{1}\partial u_{1}}=0,$
and similar for $f_{2},f_{3}$. This implies the following form of the Lagrangian where we let ${\bf v}_{0}=(x_{0},y_{0},z_{0},u_{0})$:
$L=f_{00}({\bf v}_{0},x_{1},y_{1},z_{1})+f_{01}({\bf v}_{0},x_{1},y_{1},z_{1})u_{1}+f_{10}({\bf v}_{0},x_{1},y_{1},z_{1})x_{2}+\frac{f_{11}({\bf v}_{0},y_{1},z_{1})u_{1}x_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}\\ +f_{20}({\bf v}_{0},x_{1},y_{1},z_{1})y_{2}+\frac{f_{21}({\bf v}_{0},x_{1},z_{1})u_{1}y_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}\\ +f_{30}({\bf v}_{0},x_{1},y_{1},z_{1})z_{2}+\frac{f_{31}({\bf v}_{0},x_{1},y_{1})u_{1}z_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}$
Next, splitting $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ by $x^{k}_{2}$ and $u_{1}$ we obtain conditions, among which the following are the simplest:
$\frac{\partial f_{01}}{\partial x_{1}}=\frac{\partial f_{10}}{\partial u_{0}},\quad\frac{\partial f_{01}}{\partial y_{1}}=\frac{\partial f_{20}}{\partial u_{0}},\quad\frac{\partial f_{01}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial u_{0}},\quad\frac{\partial f_{10}}{\partial y_{1}}=\frac{\partial f_{20}}{\partial x_{1}},\quad\frac{\partial f_{10}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial x_{1}},\quad\frac{\partial f_{20}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial y_{1}},$
$\displaystyle\frac{\partial f_{11}}{\partial u_{0}}=\frac{\partial f_{21}}{\partial u_{0}}=\frac{\partial f_{31}}{\partial u_{0}}=0,\quad\frac{\partial f_{11}}{\partial y_{1}}+\frac{\partial f_{21}}{\partial x_{1}}=0,\quad\frac{\partial f_{11}}{\partial z_{1}}+\frac{\partial f_{31}}{\partial x_{1}}=0,\quad\frac{\partial f_{21}}{\partial z_{1}}+\frac{\partial f_{31}}{\partial y_{1}}=0,$
@@
$L=f_{00}({\bf v}_{0},x_{1},y_{1},z_{1})+\partial_{u_{0}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})u_{1}+\partial_{x_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})x_{2}\\ +\partial_{y_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})y_{2}+\partial_{z_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})z_{2}+\frac{h_{3}(x_{0},y_{0},z_{0})y_{1}-h_{2}(x_{0},y_{0},z_{0})z_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}x_{2}\\ +\frac{h_{1}(x_{0},y_{0},z_{0})z_{1}-h_{3}(x_{0},y_{0},z_{0})x_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}y_{2}+\frac{h_{2}(x_{0},y_{0},z_{0})x_{1}-h_{1}(x_{0},y_{0},z_{0})y_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}z_{2}$
Changing this by divergence $L\mapsto L-\frac{d}{dt}\psi$ we simplify the Lagrangian as follows:
$L=h_{0}({\bf v}_{0},x_{1},y_{1},z_{1})+\frac{h_{3}(x_{0},y_{0},z_{0})y_{1}-h_{2}(x_{0},y_{0},z_{0})z_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}x_{2}\\ +\frac{h_{1}(x_{0},y_{0},z_{0})z_{1}-h_{3}(x_{0},y_{0},z_{0})x_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}y_{2}+\frac{h_{2}(x_{0},y_{0},z_{0})x_{1}-h_{1}(x_{0},y_{0},z_{0})y_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}u_{1}z_{2}$
Now again splitting $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ by $x^{k}_{2}$ and $u_{1}$ we observe the following condition
@@
Solutions of the parametrized conformal circles equation in any dimensions are obtained by rotating the motion plane to the first coordinate plane, where we get rational projective parametrization
$x(t)=p+\frac{a(t-r)+bc}{(t-r)^{2}+c^{2}},\quad y(t)=q+\frac{ac-b(t-r)}{(t-r)^{2}+c^{2}}.$
Another parametrization (in the flat space) is given by the equation (see [11] for the relation)
@@
$x(t)=a+c^{-1}\cos(ct+d),\quad y(t)=b+c^{-1}\sin(ct+d).$
@@
Now consider un-parametrized conformal geodesic equation (1) in 4D. For this the Lagrangian form $L\,dt$ should be reparametrization invariant, so we choose the coordinate $u$ as local time, and $x^{1}=x$, $x^{2}=y$, $x^{3}=z$ will depend on it. Thus $L=L(x_{i},y_{i},z_{i})$ where the derivatives are taken wrt $u$. As before, the Lagrangian can be chosen of order 22 22, i.e. $0\leq i\leq 2$ in the expression for $L$, and it should be quasi-linear, i.e. $L=f_{0}+\sum_{k=1}^{3}f_{k}x^{k}_{2}$, where $f_{i}=f_{i}(x_{0},y_{0},z_{0},x_{1},y_{1},z_{1})$. This implies for $1\leq j,k\leq 3$
@@
Thus the $3\times 3$ symbol matrix of the Euler-Lagrange equation is skew-symmetric and hence degenerate. Hence it cannot reproduce the un-parametrized conformal geodesic equation.
@@
[…]
@@
Choosing $x^{0}=t$ as parameter we get the system (again indices are raised/lowered via the Euclidean metric)
@@
Applying the same method as for parametrized conformal geodesics in 4D, one can verify that this equation is also non-variational. These tedeous computations are fully similar, so we omit the details.

### v2

@@ 1. Formulation of the Problem and the Result
Recently it was proven independently in [18] and [13] that in 3D the equation for un-parametrized conformal geodesics is variational; the corresponding Lagrangian is the torsion form $\tau\,ds$. There exists a canonical family of parametrizations of conformal geodesics, so-called projective parameter [3]. Whatever parameter is used, the equation of parametrized conformal geodesics in 3D is however not variational [13].
The higher dimensional case has remained open. One could hope that a higher dimensional version of the torsion functional (or any other conformally invariant Lagrangian) could solve the problem. That was verified in negative in [14] in the flat 4D case, using an invariant version of the variational bicomplex. The purpose of this note is to demonstrate the negative answer to this problem in full generality.
@@
Let $(M^{n},[g])$ be a conformal manifold. If its dimension $n>3$, then the equation for conformal geodesics is not variational in the classical sense neither in parametrized nor in unparametrized form.
@@ 2. Warm-up: variational calculus and circles in three and two dimensions
Let us consider curves in the Euclidean space ${\mathbb{R}}^{3}(x,y,z)$. Unparametrized (generic) curves are given by equation of the type $y=y(x)$, $z=z(x)$, where $x$ can be seen as (technical) parameter, but which can be reparametrized. The curves have natural lifts to the jet-space $J^{3}({\mathbb{R}},{\mathbb{R}}^{2})={\mathbb{R}}^{9}(x,y_{0},z_{0},y_{1},z_{1},y_{2},z_{2},y_{3},z_{3})$ and the equation for conformal geodesics (circles) determines a submanifold in this space
@@
For a Lagrangian $L=L(x,y_{0},z_{0},y_{1},z_{1},y_{2},z_{2},\dots)$ the Euler ($y$- and $z$-) operators are
$\frac{\delta L}{\delta y}=\frac{\partial L}{\partial y_{0}}-\frac{d}{dx}\Bigl(\frac{\partial L}{\partial y_{1}}\Bigr)+\frac{d^{2}}{dx^{2}}\Bigl(\frac{\partial L}{\partial y_{2}}\Bigr)-\dots,\qquad\frac{\delta L}{\delta z}=\frac{\partial L}{\partial z_{0}}-\frac{d}{dx}\Bigl(\frac{\partial L}{\partial z_{1}}\Bigr)+\frac{d^{2}}{dx^{2}}\Bigl(\frac{\partial L}{\partial z_{2}}\Bigr)-\dots.$ (4)
The Euler-Lagrange equations EL are
@@
The inverse variational problem is to detect if the given equation, in our case (3), is of form (5) for some $L$. Let us stress that we do not require that the first equation of (3) is identical with the first Euler operator in (4), etc, but the equations have to coincide as submanifolds in jets (inverse problem for differential operators is simpler and is solved via the so-called Helmholtz criterion [2]).
@@
The Lagrangian for variational equation of order $k$ can be chosen to be of order $<k$: this is achieved by modification of $L$ via divergence terms (integration by parts) and is known as Vainberg–Tonti method, see e.g. [15]. If $L$ is of order 2 then the coefficients of $y_{4},z_{4}$ in EL are $\frac{\partial^{2}L}{\partial y_{2}^{2}}$, $\frac{\partial^{2}L}{\partial y_{2}\partial z_{2}}$, $\frac{\partial^{2}L}{\partial z_{2}^{2}}$. Since (3) is of the third order, these expressions have to vanish and we get
$L=f(y_{0},z_{0},y_{1},z_{1})+h(y_{0},z_{0},y_{1},z_{1})y_{2}+r(y_{0},z_{0},y_{1},z_{1})z_{2}.$
Using the fact, that the Lagrangian is defined up to divergence, i.e. that the transformation $L\mapsto L+\frac{d}{dx}f$ does not change the Euler-Lagrange equations, we can eliminate the coefficient $r\mapsto 0$. Then we set our equations (3) into EL and split it by 2-jets $y_{2},z_{2}$, receiving the following overdetermined PDE system:
$\displaystyle y_{1}^{2}h_{y_{0},y_{0}}+2y_{1}z_{1}h_{y_{0},z_{0}}+z_{1}^{2}h_{z_{0},z_{0}}=y_{1}f_{y_{0},y_{1}}+z_{1}f_{z_{0},y_{1}}-f_{y_{0}},\ y_{1}f_{y_{0},z_{1}}+z_{1}f_{z_{0},z_{1}}=f_{z_{0}},$
$\displaystyle f_{y_{1},z_{1}}=h_{z_{0}},\ f_{z_{1},z_{1}}=0,\ y_{1}h_{y_{0},z_{1}}+z_{1}h_{z_{0},z_{1}}=0,\ y_{1}h_{y_{0},y_{1}}+z_{1}h_{z_{0},y_{1}}=f_{y_{1},y_{1}}-2h_{y_{0}},$
$\displaystyle(1+y_{1}^{2}+z_{1}^{2})h_{y_{1},z_{1}}=-3y_{1}h_{z_{1}},\ (1+y_{1}^{2}+z_{1}^{2})h_{z_{1},z_{1}}=-3z_{1}h_{z_{1}}.$
@@
This Lagrangian slightly simplifies the one found in formula (11) of [13]. Prior to this a third order Lagrangian for conformal circles in Euclidean space was established in [4]. Note that only this latter Lagrangian (torsion) is reparametrization-invariant. The others have this property modulo divergence.
@@
Considering the equation for parametrized circles (see also Remark 1 later)
$x_{3}=-(x_{2}^{2}+y_{2}^{2})x_{1},\quad y_{3}=-(x_{2}^{2}+y_{2}^{2})y_{1},$
@@
$x_{3}=3\frac{x_{1}x_{2}+y_{1}y_{2}}{x_{1}^{2}+y_{1}^{2}}x_{2}-\frac{3}{2}\frac{x_{2}^{2}+y_{2}^{2}}{x_{1}^{2}+y_{1}^{2}}x_{1},\quad y_{3}=3\frac{x_{1}x_{2}+y_{1}y_{2}}{x_{1}^{2}+y_{1}^{2}}y_{2}-\frac{3}{2}\frac{x_{2}^{2}+y_{2}^{2}}{x_{1}^{2}+y_{1}^{2}}y_{1}.$
We start with the Lagrangian $L=L(t,x_{0},y_{0},x_{1},y_{1},x_{2},y_{2})$ and then the condition that EL is of order $<4$ and simplification modulo divergences implies that we can take $L=f(t,x_{0},y_{0},x_{1},y_{1})+h(t,x_{0},y_{0},x_{1},y_{1})y_{2}$. However one of the EL equations is $h_{x_{1}}=0$, so $h=h(t,x_{0},y_{0},y_{1})$ and hence a change by a divergences can bring the Lagrangian to the form $L=f(t,x_{0},y_{0},x_{1},y_{1})$, for which the Euler-Lagrange equation is of order at most 2. This proves that in dimension 2 the analog equation for (conformal) circles (both in parametric and unparametric form) is not variational.
@@ 3. The inverse problem for conformal circles in four and five dimensions
Using the Vainberg–Tonti method, we can choose $L=L\bigl(x_{0}^{j},x_{1}^{j},x_{2}^{j}:1\leq j\leq 4\bigr)$; the general Lagrangian may also depend on $t$ but this does not change the argument below and so will be omitted. Denote the Euler-Lagrange operator
@@
$f_{1;3}f_{2;4}-f_{1;4}f_{2;3}+f_{1;4}f_{3;2}-f_{1;2}f_{3;4}+f_{1;2}f_{4;3}-f_{1;3}f_{4;2}\\ +f_{2;1}f_{3;4}-f_{2;4}f_{3;1}+f_{2;3}f_{4;1}-f_{2;1}f_{4;3}+f_{3;1}f_{4;2}-f_{3;2}f_{4;1}$ (10)
which thus should be nonzero. Here $f_{i;j}$ denotes the partial derivative of $f_{i}$ by $x^{j}_{1}$. If we freeze 0-jet variables $x^{j}_{0}$ getting $df_{i}=\sum_{j=1}^{4}f_{i;j}dx^{j}_{1}$, then the nonvanishing of expression (10) is equivalent to nonvanishing of the following 4-form, where the indices $ijkl$ in the summation run over even permutations of $1234$:
$\sum df_{i}\wedge df_{j}\wedge dx_{1}^{k}\wedge dx_{1}^{l}\neq 0.$ (11)
@@
$\displaystyle\frac{\partial^{2}f_{0}}{\partial w_{1}\partial w_{1}}=0,\quad 3w_{1}\frac{\partial f_{1}}{\partial w_{1}}+(x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+w_{1}^{2})\frac{\partial^{2}f_{1}}{\partial w_{1}\partial w_{1}}=0,\quad 3x_{1}\frac{\partial f_{1}}{\partial w_{1}}+(x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+w_{1}^{2})\frac{\partial^{2}f_{1}}{\partial x_{1}\partial w_{1}}=0,$
and similarly for $f_{2},f_{3}$. This implies the following form of the Lagrangian where we let ${\bf v}_{0}=(x_{0},y_{0},z_{0},w_{0})$:
$L=f_{00}({\bf v}_{0},x_{1},y_{1},z_{1})+f_{01}({\bf v}_{0},x_{1},y_{1},z_{1})w_{1}+f_{10}({\bf v}_{0},x_{1},y_{1},z_{1})x_{2}+\frac{f_{11}({\bf v}_{0},y_{1},z_{1})w_{1}x_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+w_{1}^{2}}}\\ +f_{20}({\bf v}_{0},x_{1},y_{1},z_{1})y_{2}+\frac{f_{21}({\bf v}_{0},x_{1},z_{1})w_{1}y_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+w_{1}^{2}}}\\ +f_{30}({\bf v}_{0},x_{1},y_{1},z_{1})z_{2}+\frac{f_{31}({\bf v}_{0},x_{1},y_{1})w_{1}z_{2}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+w_{1}^{2}}}.$
Next, splitting $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ by $x^{k}_{2}$ and $w_{1}$ we obtain conditions, among which the following are the simplest:
$\frac{\partial f_{01}}{\partial x_{1}}=\frac{\partial f_{10}}{\partial w_{0}},\quad\frac{\partial f_{01}}{\partial y_{1}}=\frac{\partial f_{20}}{\partial w_{0}},\quad\frac{\partial f_{01}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial w_{0}},\quad\frac{\partial f_{10}}{\partial y_{1}}=\frac{\partial f_{20}}{\partial x_{1}},\quad\frac{\partial f_{10}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial x_{1}},\quad\frac{\partial f_{20}}{\partial z_{1}}=\frac{\partial f_{30}}{\partial y_{1}},$
$\displaystyle\frac{\partial f_{11}}{\partial w_{0}}=\frac{\partial f_{21}}{\partial w_{0}}=\frac{\partial f_{31}}{\partial w_{0}}=0,\quad\frac{\partial f_{11}}{\partial y_{1}}+\frac{\partial f_{21}}{\partial x_{1}}=0,\quad\frac{\partial f_{11}}{\partial z_{1}}+\frac{\partial f_{31}}{\partial x_{1}}=0,\quad\frac{\partial f_{21}}{\partial z_{1}}+\frac{\partial f_{31}}{\partial y_{1}}=0,$
@@
$L=f_{00}({\bf v}_{0},x_{1},y_{1},z_{1})+\partial_{w_{0}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})w_{1}+\partial_{x_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})x_{2}\\ +\partial_{y_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})y_{2}+\partial_{z_{1}}\psi({\bf v}_{0},x_{1},y_{1},z_{1})z_{2}+\frac{h_{3}(x_{0},y_{0},z_{0})y_{1}-h_{2}(x_{0},y_{0},z_{0})z_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}x_{2}\\ +\frac{h_{1}(x_{0},y_{0},z_{0})z_{1}-h_{3}(x_{0},y_{0},z_{0})x_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}y_{2}+\frac{h_{2}(x_{0},y_{0},z_{0})x_{1}-h_{1}(x_{0},y_{0},z_{0})y_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}z_{2}.$
Changing this by divergence $L\mapsto L-\frac{d}{dt}\psi$ we once more simplify the Lagrangian as follows:
$L=h_{0}({\bf v}_{0},x_{1},y_{1},z_{1})+\frac{h_{3}(x_{0},y_{0},z_{0})y_{1}-h_{2}(x_{0},y_{0},z_{0})z_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}x_{2}\\ +\frac{h_{1}(x_{0},y_{0},z_{0})z_{1}-h_{3}(x_{0},y_{0},z_{0})x_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}y_{2}+\frac{h_{2}(x_{0},y_{0},z_{0})x_{1}-h_{1}(x_{0},y_{0},z_{0})y_{1}}{(x_{1}^{2}+y_{1}^{2}+z_{1}^{2})\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}+u_{1}^{2}}}w_{1}z_{2}.$
Now again splitting $\mathop{\rm ev}\nolimits_{\mathcal{E}}EL_{j}$ by $x^{k}_{2}$ and $w_{1}$ we observe the following condition
@@
Solutions of the parametrized conformal circles equation in any dimensions (1) are obtained by rotating the motion plane to the first coordinate plane, where we get rational projective parametrization
$x(t)=p+\frac{\alpha(t-r)+\beta\gamma}{(t-r)^{2}+\gamma^{2}},\quad y(t)=q+\frac{\alpha\gamma-\beta(t-r)}{(t-r)^{2}+\gamma^{2}}.$
Note that this projective parameter is the solution to a third order ODE: the scalar product of (1) with $u$ gives $(u,\dot{a})=3|u|^{-2}(u,a)^{2}-\tfrac{3}{2}|a|^{2}$.
Another parametrization (in the conformally flat space given by the metric of nonzero constant curvature; see e.g. [13]) is given by the equation
@@
$x(t)=\alpha+\gamma^{-1}\cos\bigl(\gamma(t-r)\bigr),\quad y(t)=\beta+\gamma^{-1}\sin\bigl(\gamma(t-r)\bigr).$ (12)
@@
Now consider un-parametrized conformal geodesic equation (2) in 4D. For this the Lagrangian form $L\,dt$ should be reparametrization invariant, so we choose the coordinate $w$ as local time, and $x^{1}=x$, $x^{2}=y$, $x^{3}=z$ will depend on it. Thus $L=L(x_{i},y_{i},z_{i})$ where the derivatives are taken wrt $w$. As before, the Lagrangian can be chosen of order $2$, i.e. $0\leq i\leq 2$ in the expression for $L$, and it should be quasi-linear, i.e. $L=f_{0}+\sum_{k=1}^{3}f_{k}x^{k}_{2}$, where $f_{i}=f_{i}(x_{0},y_{0},z_{0},x_{1},y_{1},z_{1})$. This implies for $1\leq j,k\leq 3$
@@
Thus the $3\times 3$ symbol matrix of the Euler-Lagrange equation is skew-symmetric and hence degenerate. Therefore it cannot reproduce the un-parametrized conformal geodesic equation.
@@
In 5D the equation for parametrized conformal geodesics is trivially non-variational, as follows from the above remark. Un-parametrized conformal geodesics are given by the following system (where the bracket $[\cdot\cdot]$ stands for skew-symmetrization over the indicated indices)
@@
Choosing $x^{0}=t$ as parameter we get the system (the indices are lowered via the Euclidean metric)
@@
Applying the same method as for parametrized conformal geodesics in 4D, one can verify that this equation is also non-variational. These tedious computations are fully similar, so we omit the details.